OpenStudy (anonymous):

If the area of a rectangle with width x can be represented with the expression A(x) = x(14 – x), what is the perimeter of the rectangle?

OpenStudy (jango_in_dtown):

area=length * width, given A=x(14-x) since x is width, so length is 14-x perimeter=2(length+width)=2(14-x+x)=2*14=28

OpenStudy (jango_in_dtown):

@undefeated2015